A 32-inch board is to be cut into three pieces so that the second piece is three times as long as the first piece and the third piece is 44 times as long as the first piece. If x represents the length of the first piece, find the lengths of all three pieces.
Answer by Cromlix(4381)

Hi there,
First piece = x
Second piece = 3x
Third piece = 44x
x + 3x + 44x = 32 inches
48x = 32 inches
x = 32/48 inches
x = 2/3 inches
First piece = 2/3 of an inch.
Second piece = 2 inches.
Third piece = 29 1/3 inches.
Hope this helps :-)